OpenStudy (anonymous):

Simplify the following expression. (7^5)^3 × 7^7 A. 7^8 B. 7^15 C. 7^56 D. 7^22

OpenStudy (jdoe0001):

\(\bf (7^5)^3 \times 7^7\implies 7^{{\color{red}{ 5\cdot 3}}}\times 7^{\color{blue}{ 7}}\implies 7^{{\color{red}{ 5\cdot 3}}+{\color{blue}{ 7}}}\)

OpenStudy (anonymous):

@jdoe0001 i don't understand??

jimthompson5910 (jim_thompson5910):

First simplify (7^5)^3

jimthompson5910 (jim_thompson5910):

We use the rule (a^b)^c = a^(b*c) ... in other words, multiply the exponents

jimthompson5910 (jim_thompson5910):

so (7^5)^3 = 7^(5*3) = 7^(15)

jimthompson5910 (jim_thompson5910):

That turns (7^5)^3 × 7^7 into 7^(15) × 7^7

jimthompson5910 (jim_thompson5910):

Now use the rule a^b * a^c = a^(b+c) ... in other words, add the exponents

OpenStudy (anonymous):

so it's 7^22?

jimthompson5910 (jim_thompson5910):

7^(15) × 7^7 = 7^(15 + 7) = 7^(22) So that means (7^5)^3 × 7^7 = 7^(22)

OpenStudy (anonymous):

@jim_thompson5910 yay thank you so much for explaining it out for me it makes much more sense now

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

its D

OpenStudy (anonymous):

@samantha6233 i know thanks